# Representations of categories of finite relational structures and associated endomorphism monoids

### 18 Mar 2026 v1

@@ 3. Highly homogeneous structures and mutation graphs
Recall that the action of $G$ on $\Omega$ is highly homogeneous if, for every $n\in\mathbb{Z}_{+}$, the set of $n$-element subsets of $\Omega$ forms a single $G$-orbit. By the fundamental classification theorem of Cameron, there are, up to equivalence, exactly five highly homogeneous relational structures on $\Omega$ induced by such group actions: the empty relation (pure set), the dense linear order, the dense cyclic order, the dense linear betweenness relation, and the dense separation relation; see [2, Section 13].
Accordingly, a skeletal category $\mathscr{C}$} of finite substructures and relation-preserving maps can be described explicitly as follows:
@@
For the other relational structures, the mutation graph $\Gamma^{\prime}_{n-1,n}$ is a subgraph of $\Gamma_{n-1,n}$. Since $\Gamma_{n-1,n}$ is a bipartite graph, so is $\Gamma^{\prime}_{n-1,n}$. ∎
@@
The map $f^{\prime}$ is clearly an injective map of sets and satisfies $h\circ f^{\prime}=h\circ f$ by construction. Besides, $f^{\prime}$ is a morphism in $\mathscr{C}$}. To see this, note that the relational structure on $h(f([m]))\subseteq[n-1]$ is induced by the relations on $[n]$ via the morphism $h$. Since $h\circ f^{\prime}=h\circ f$ and $h$ is a morphism that identifies only the pair $\{a,b\}$, the fact that $h\circ f$ is an embedding implies that the relations on the tuple $(f^{\prime}(1),\dots,f^{\prime}(m))$ must coincide with those on $(f(1),\dots,f(m))$.
@@ 4. Normalization
We can deduce several results from the above equivalence. In the rest of this section let $\mathscr{C}$} be one of the five categories associated to highly homogeneous relational structures (including $\mathrm{FA}$}).
@@ 5. Standard modules
In the rest of this section let $k$ be a field of characteristic 0, although many results still hold for $\mathrm{OA}$} for any field, and for $\mathrm{BA}$} for fields with characteristic different from 2. For a fixed object $[n]$ in $\mathscr{C}$}, the group algebra $kG_{n}$ is semisimple. Denote by $\operatorname{Irr}(G_{n})$ the set of irreducible representations (up to isomorphism) of $G_{n}$. For a fixed $\lambda\in\operatorname{Irr}(G_{n})$, recall
  • $P_{n,\lambda}=P_{n}\otimes_{kG_{n}}\lambda$;
  • $\mathfrak{I}_{n,\lambda}=\mathfrak{I}_{n}\otimes_{kG_{n}}\lambda$;
  • $\Delta_{n,\lambda}=P_{n,\lambda}/\mathfrak{I}_{n,\lambda}\cong\Delta_{n}\otimes_{kG_{n}}\lambda$.
Call $\Delta_{n,\lambda}$ an indecomposable standard module as it is an indecomposable direct summand of $\Delta_{n}$.
@@
    such that $I^{n}/I^{n-1}\cong\Delta_{n,\lambda}$ and $I^{i}/I^{i-1}$ is a finite coproduct of some $\Delta_{i,\mu}$;
@@
Statement (1) follows from the proofs of [9, Lemma 3.10 and Proposition 3.11], and the last two statements follow from [9, Theorem 3.13] and its proof. ∎
@@
If there is a nonzero homomorphism from $\Delta_{n+1,\lambda}$ to $\Delta_{n,\mu}$, then $(\lambda,\mu)$ must be the pair of irreducible representations appearing in the following table.
@@
Let $k$ be a field of characteristic 0. Then every regular indecomposable standard module is irreducible.
@@
Suppose, for contradiction, that there exists a regular indecomposable standard module $\Delta_{n,\lambda}$ which is not irreducible. By Corollary 4.9, $\Delta_{n,\lambda}$ has a nonzero socle. Take an irreducible $\mathscr{C}$}-module $L_{m,\mu}$ in this socle. Then $L_{m,\mu}$ is a proper submodule of $\Delta_{n,\lambda}$. Besides, we get a nonzero homomorphism $f:\Delta_{m,\mu}\to\Delta_{n,\lambda}$ whose image is precisely $L_{m,\mu}$.
By Theorem 5.1, this can occur if and only if $m=n$ or $m=n+1$. But clearly $m\neq n$. Otherwise we obtain a nonzero $kG_{n}$-module homomorphism
$f_{n}:\Delta_{m,\mu}(n)\cong\mu\longrightarrow\Delta_{n,\lambda}(n)\cong\lambda$
which must be an isomorphism since $\mu$ and $\lambda$ are simple $kG_{n}$-modules. Consequently, $f$ is surjective since $\Delta_{n,\lambda}$ is generated by $\lambda$, contradicting the assumption that $L_{m,\mu}$ is a proper submodule of $\Delta_{n,\lambda}$. Thus $m=n+1$. But by Proposition 5.4, this implies that $\Delta_{n,\lambda}$ is singular, contradicting the hypothesis. Therefore $\Delta_{n,\lambda}$ must be irreducible. ∎
Now we consider singular indecomposable standard modules.
Proposition 5.6.
Let $k$ be a field of characteristic 0. Then every singular indecomposable standard module has length 2.
@@
Let $\Delta_{n,\lambda}$ be a singular indecomposable standard module, and let $\mathscr{D}$} be the simplex category, viewed as a subcategory of $\mathscr{C}$}. As in the proof of Corollary 4.8, we have
@@
$\operatorname{Res}^{{\mathscr{C}}}_{{\mathscr{D}}}\Delta_{n,\lambda}=\operatorname{Res}_{{\mathscr{D}}}^{{\mathscr{C}}}(\Delta_{n}\otimes_{kG_{n}}\lambda)\cong(\Delta_{n}^{{\mathscr{D}}}\otimes_{k}kG_{n})\otimes_{kG_{n}}\lambda\cong\Delta_{n}^{{\mathscr{D}}}\otimes_{k}\lambda\cong\Delta_{n}^{{\mathscr{D}}},$
since $\lambda$ is one-dimensional.
Note that $\Delta_{n}^{{\mathscr{D}}}$ is the normalized projective $\mathscr{D}$}-module. By the Dold-Kan correspondence, it has a simple top and a simple socle, and no further composition factors. Hence its length is $2$. Since the restriction functor is exact and cannot send a nonzero module to zero (as $\mathscr{D}$} is a wide subcategory of $\mathscr{C}$}), it follows that $\Delta_{n,\lambda}$ has length at most $2$ as a $\mathscr{C}$}-module. Furthermore, since $\Delta_{n,\lambda}$ is singular, there exists a nonzero homomorphism $\phi:\Delta_{n+1,\mu}\to\Delta_{n,\lambda}$. Clearly, $\phi$ is not surjective since the value of the first module on $[n]$ is zero while the value of the second module is nonzero. Consequently, $\Delta_{n,\lambda}$ has length at least $2$, which forces its length to be exactly $2$. ∎
@@
Let $k$ be a field of characteristic 0. Then the isomorphism classes of irreducible $\mathscr{C}$}-modules are parameterized by elements in the set
@@
    if $\lambda$ is regular, then the corresponding irreducible $\mathscr{C}$}-module is isomorphic to $\Delta_{n,\lambda}$;
@@
    if $\lambda$ is singular, then the corresponding irreducible $\mathscr{C}$}-module is the top of $\Delta_{n,\lambda}$ or the socle of $\Delta_{n+1,\mu}$ (for $n\geqslant 2$), where $(\lambda,\mu)$ is a pair of singular irreducible representations appearing in Proposition 5.4.
@@
The Grothendieck group $K_{0}({\mathscr{C}})$ of the category of finitely generated $\mathscr{C}$}-modules has the following canonical decomposition:
@@
Next, we show that the set of isomorphism classes of indecomposable standard modules form a basis as well. By Proposition 5.5 and 5.6, if $\lambda\in\operatorname{Irr}(G_{n})$ is regular, then $\Delta_{n,\lambda}]=[L_{n,\lambda}]$; otherwise,
@@
for a unique singular $\mu\in\operatorname{Irr}(G_{n+1})$. In both cases, the class $\Delta_{n,\lambda}]$ is equal to $L_{n,\lambda}]$ plus (possibly) a simple module of strictly higher degree. Therefore, the transition matrix between $\{[L_{n,\lambda}]\}$ and $\{[\Delta_{n,\lambda}]\}$ is upper unitriangular. Since an upper unitriangular matrix (even of infinite size, as long as the degrees are bounded below) is invertible over $\mathbb{Z}$, the classes $\{[\Delta_{n,\lambda}]\}$ form an alternative basis for $K_{0}({\mathscr{C}})$. ∎
@@ 6. Further results for $\mathrm{CA}$}, $\mathrm{SA}$} and $\mathrm{FA}$}
In the case that $k$ is a field of characteristic 0, the representation theory of $\mathrm{OA}$} and $\mathrm{BA}$} is very transparent. Thus in this section we only consider representations of the other three categories.
@@ 6.1. $\mathrm{CA}$} and $\mathrm{SA}$}
In the rest of this section let $k$ be a field of characteristic 0. Since $d$ is either 1 or $2$, this computation tells us that most indecomposable standard modules are actually projective. More precisely, fix $n\geqslant 1$, and let
@@
$k^{d}\cong\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n})=\bigoplus_{i=1}^{r}\operatorname{Ext}_{k{\mathscr{C}}}^{1}((\Delta_{n+1,\mu_{i}})^{\dim_{k}\mu_{i}},\,\Delta_{n})$
@@
$\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n})=\bigoplus_{i=1}^{r}\bigoplus_{j=1}^{s}\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1,\mu_{j}},\,\Delta_{n,\lambda_{i}})$
@@
$\Delta_{n,\lambda}(n+1)\cong k{\mathscr{C}}^{+}(n,n+1)\otimes_{kG_{n}}\lambda,$
which has dimension $n+1$ since $k{\mathscr{C}}^{+}(n,n+1)$ is a right free $kG_{n}$-module of rank $n+1$. For the third term, we note that $\mathfrak{I}_{n+1}$ has a short exact sequence
@@
has dimension $n+1$. Consequently, $\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n,\lambda})$ has dimension 1, so it is nonzero as claimed.
Let $e_{\mu}\in kG_{n+1}$ be a central idempotent such that $kG_{n+1}e_{\mu}\cong\mu$. Applying $\operatorname{Hom}_{k{\mathscr{C}}}(-,\,\Delta_{n})$ to the short exact sequence
@@
which has dimension $n$ since $\Delta_{n}(n+1)\cong k{\mathscr{C}}^{+}(n,n+1)$ is a free left $kG_{n+1}$-module of rank $n$. Finally, $\mathfrak{I}_{n+1}e_{\mu}$ admits a filtration whose only contribution to $\operatorname{Hom}(-,\,\Delta_{n})$ comes from
$k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}k{\mathscr{C}}^{-}(n+1,n)e_{\mu}\cong k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}(kG_{n}\otimes_{k}kG_{n+1}e_{\mu})\cong k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}(kG_{n}\otimes_{k}\mu)\cong\Delta_{n},$
@@
$\operatorname{Hom}_{k{\mathscr{C}}}(\mathfrak{I}_{n+1}e_{\mu},\,\Delta_{n})\cong\operatorname{Hom}_{k{\mathscr{C}}}(\Delta_{n},\,\Delta_{n})\cong kG_{n}$
has dimension $n$, so $\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1,\mu},\,\Delta_{n})$ has dimension 1, which is nonzero as claimed.
For the category $\mathrm{SA}$}, the argument proceeds similarly with a refined look at the bimodule structure of the morphisms. For $n\geqslant 3$, $G_{n+1}=D_{2(n+1)}$ and $G_{n}=D_{2n}$ are dihedral, so $k{\mathscr{C}}^{-}(n+1,n)$ is not a free bimodule. Instead, one has
@@
$k{\mathscr{C}}^{-}(n+1,n)e_{\mu}\cong kG_{n}\otimes_{kC_{2}}\mu|_{C_{2}}.$
Since $\mu$ is a one-dimensional character ($\lambda^{\pm}$) whose restriction to $C_{2}$ remains one-dimensional, the resulting module is induced from a one-dimensional representation of $C_{2}$ to $G_{n}$, which has dimension $n$. Consequently, the dimension of $\operatorname{Hom}_{k{\mathscr{C}}}(\mathfrak{I}_{n+1}e_{\mu},\,\Delta_{n})$ remains $n$, preserving the exact dimension balance required to show that $\operatorname{Ext}^{1}_{k{\mathscr{C}}}(\Delta_{n+1,\mu},\,\Delta_{n,\lambda})$ is one-dimensional. ∎

### 25 Apr 2026 v3

@@ 3. Highly homogeneous structures and mutation graphs
Recall that the action of a permutation group $G\leqslant\operatorname{Sym}(\Omega)$ on $\Omega$ is highly homogeneous if, for every $n\in\mathbb{Z}_{+}$, the set of $n$-element subsets of $\Omega$ forms a single $G$-orbit. By the fundamental classification theorem of Cameron, there are, up to equivalence, exactly five highly homogeneous relations $\mathcal{R}_{\Omega}$ on $\Omega$ induced by such group actions: the empty relation (pure set), the dense linear order, the dense cyclic order, the dense linear betweenness relation, and the dense separation relation; see [2, Section 13].
Accordingly, a skeletal category $\mathscr{C}$} of finite subsets and relation-preserving maps can be described explicitly as follows:
@@
For the other highly homogeneous relations, the mutation graph $\Gamma^{\prime}_{n-1,n}$ is a subgraph of $\Gamma_{n-1,n}$. Since $\Gamma_{n-1,n}$ is a bipartite graph, so is $\Gamma^{\prime}_{n-1,n}$. ∎
@@
The map $f^{\prime}$ is clearly an injective map of sets and satisfies $h\circ f^{\prime}=h\circ f$ by construction. Besides, $f^{\prime}$ is a morphism in $\mathscr{C}$}. To see this, note that the relation on $h(f([m]))\subseteq[n-1]$ is induced by the relations on $[n]$ via the morphism $h$. Since $h\circ f^{\prime}=h\circ f$ and $h$ is a morphism that identifies only the pair $\{a,b\}$, the fact that $h\circ f$ is an embedding implies that the relations on the tuple $(f^{\prime}(1),\dots,f^{\prime}(m))$ must coincide with those on $(f(1),\dots,f(m))$.
@@ 4. Normalization
We can deduce several results from the above equivalence. In the rest of this section let $\mathscr{C}$} be one of the five categories associated to highly homogeneous relations (including $\mathrm{FA}$}).
@@ 5. Standard modules
In the rest of this section let $k$ be a field. For a fixed object $[n]$ in $\mathscr{C}$}, denote by $\operatorname{Irr}(G_{n})$ the set of irreducible representations (up to isomorphism) of $G_{n}$. For a fixed $\lambda\in\operatorname{Irr}(G_{n})$, let $e_{\lambda}\in kG_{n}$ be the primitive idempotent such that $kG_{n}e_{\lambda}$ is the projective cover of $\lambda$. Define
  • $P_{n,\lambda}=P_{n}\otimes_{kG_{n}}kG_{n}e_{\lambda}$;
  • $\mathfrak{I}_{n,\lambda}=\mathfrak{I}_{n}\otimes_{kG_{n}}kG_{n}e_{\lambda}$;
  • $\Delta_{n,\lambda}=P_{n,\lambda}/\mathfrak{I}_{n,\lambda}\cong\Delta_{n}\otimes_{kG_{n}}kG_{n}e_{\lambda}$.
Here the isomorphism holds since $\Delta_{n}$ is a right free $kG_{n}$-module, so the functor $\Delta_{n}\otimes_{kG_{n}}$- is exact. Indeed, for each $l\geqslant 1$, the value $\Delta_{n}(l)$ of $\Delta_{n}$ on $[l]$ is isomorphic to $k{\mathscr{C}}^{+}(n,l)$, which is a right free $kG_{n}$-module since $G_{n}$ acts freely on ${\mathscr{C}}^{+}(n,l)$ from the right side. Call $\Delta_{n,\lambda}$ an indecomposable standard module as it is an indecomposable direct summand of $\Delta_{n}$ corresponding the primitive idempotent $e_{\lambda}$.
@@
    such that $I^{n}/I^{n-1}\cong\Delta_{n,\lambda}$ and $I^{i}/I^{i-1}$ is a finite coproduct of some $\Delta_{i,\mu}$ with $i<n$;
@@
Statement (1) follows from the proofs of [11, Lemma 3.10 and Proposition 3.11], and the last two statements follow from [11, Theorem 3.13] and its proof. ∎
@@
Let $k$ be a field with characteristic different from 2. If there is a nonzero homomorphism from $\Delta_{n+1,\lambda}$ to $\Delta_{n,\mu}$, then $(\lambda,\mu)$ must be the pair of irreducible representations appearing in the following table.
@@
Remark 5.5.
The assumption that the characteristic of $k$ is distinct from 2 guarantees that $v_{+}$ and $v_{-}$ in the above proof are linearly independent. The reader can see that this assumption is not necessary for $\mathrm{OA}$}, $\mathrm{CA}$} or $\mathrm{FA}$}.
If the characteristic of $k$ is 2, then the one-dimensional representations of the dihedral groups (or cyclic groups of order 2) coincide. More precisely:
  • for $\mathrm{SA}$}, we have $\lambda^{+}=\lambda^{-}$ and $\mu^{+}=\mu^{-}$, so there is only one singular pair $(\lambda,\mu)$;
  • for $\mathrm{BA}$}, we have $\mathrm{triv}=\operatorname{sgn}$, so there is only one singular pair $(\mathrm{triv},\mathrm{triv})$.
@@
To classify irreducible $\mathscr{C}$}-modules, given $\lambda\in\operatorname{Irr}(G_{n})$, we define
$\overline{\Delta}_{n,\lambda}=\Delta_{n}\otimes_{kG_{n}}\lambda,$
which is a quotient of $\Delta_{n,\lambda}$. If $kG_{n}$ is semisimple, then $\overline{\Delta}_{n,\lambda}=\Delta_{n,\lambda}$, but in general they are different.
@@
Let $k$ be a field. If $\lambda\in\operatorname{Irr}(G_{n})$ is regular, then $\overline{\Delta}_{n,\lambda}$ is irreducible.
@@
Suppose, for contradiction, that there exists a $\lambda\in\operatorname{Irr}(G_{n})$ such that $\overline{\Delta}_{n,\lambda}$ is not irreducible. By Corollary 4.9, $\overline{\Delta}_{n,\lambda}$ has a nonzero socle. Take an irreducible $\mathscr{C}$}-module $L_{m,\mu}$ in this socle. Then $L_{m,\mu}$ is a proper submodule of $\overline{\Delta}_{n,\lambda}$. Besides, we get a nonzero homomorphism
$f:\Delta_{m,\mu}\longrightarrow\overline{\Delta}_{n,\lambda}$
whose image is precisely $L_{m,\mu}$.
Note that the functor $\Delta_{n}\otimes_{kG_{n}}$- is exact. Also note that $\lambda$ is the socle of $kG_{n}e_{\lambda}$ since $kG_{n}$ is a symmetric algebra. Therefore, applying $\Delta_{n}\otimes_{kG_{n}}$- to the inclusion $\lambda\to kG_{n}e_{\lambda}$ we get an injective $\mathscr{C}$}-module homomorphism
$g:\overline{\Delta}_{n,\lambda}=\Delta_{n}\otimes_{kG_{n}}\lambda\longrightarrow\Delta_{n}\otimes_{kG_{n}}kG_{n}e_{\lambda}=\Delta_{n,\lambda}.$
Consequently, we obtain a nonzero $\mathscr{C}$}-module homomorphism
$g\circ f:\Delta_{m,\mu}\longrightarrow\Delta_{n,\lambda}.$
By Theorem 5.1, this can occur if and only if $m=n$ or $m=n+1$.
If $m=n$, we obtain a nonzero $kG_{n}$-module homomorphism
$f_{n}:\Delta_{n,\mu}(n)\cong kG_{n}e_{\mu}\longrightarrow\overline{\Delta}_{n,\lambda}(n)\cong\lambda.$
This happens if and only if $\mu\cong\lambda$. Accordingly, $f_{n}$ is surjective, so is $f$ since $\overline{\Delta}_{n,\lambda}$ is generated by $\lambda$, contradicting the assumption that $L_{m,\mu}$ is a proper submodule of $\overline{\Delta}_{n,\lambda}$. Thus $m=n+1$. But by Proposition 5.4, this implies that $\lambda$ is singular, contradicting the hypothesis. Therefore $\overline{\Delta}_{n,\lambda}$ must be irreducible. ∎
Now we consider the singular case.
Lemma 5.7.
Let $k$ be a field. If $\lambda\in\operatorname{Irr}(G_{n})$ is singular, then $\overline{\Delta}_{n,\lambda}$ has length at most 2.
@@
Let $\mathscr{D}$} be the simplex category, viewed as a subcategory of $\mathscr{C}$}. As in the proof of Corollary 4.8, we have
@@
$\operatorname{Res}^{{\mathscr{C}}}_{{\mathscr{D}}}\overline{\Delta}_{n,\lambda}=\operatorname{Res}_{{\mathscr{D}}}^{{\mathscr{C}}}(\Delta_{n}\otimes_{kG_{n}}\lambda)\cong(\Delta_{n}^{{\mathscr{D}}}\otimes_{k}kG_{n})\otimes_{kG_{n}}\lambda\cong\Delta_{n}^{{\mathscr{D}}}\otimes_{k}\lambda\cong\Delta_{n}^{{\mathscr{D}}},$
since the singular $\lambda$ is one-dimensional as shown in Proposition 5.4.
Note that $\Delta_{n}^{{\mathscr{D}}}$ is the normalized projective $\mathscr{D}$}-module. By the Dold-Kan correspondence, it has a simple top and a simple socle, and no further composition factors. Hence its length is $2$. Since the restriction functor is exact and cannot send a nonzero module to zero (as $\mathscr{D}$} is a wide subcategory of $\mathscr{C}$}), it follows that $\overline{\Delta}_{n,\lambda}$ has length at most $2$ as a $\mathscr{C}$}-module. ∎
Lemma 5.8.
Let $k$ be a field and $\lambda\in\operatorname{Irr}(G_{n})$ be a singular irreducible representation. Then there exists a nonzero non-surjective homomorphism
$\overline{\phi}:\overline{\Delta}_{n+1,\mu}\longrightarrow\overline{\Delta}_{n,\lambda}$
with $\mu\in\operatorname{Irr}(G_{n+1})$ singular.
Proof.
Let $\{\lambda_{i}\mid i\in[r]\}$ be the composition factors of $kG_{n}e_{\lambda}$ where $e_{\lambda}\in kG_{n}$ is a primitive idempotent such that $kG_{n}e_{\lambda}$ is the projective cover of $\lambda$. Then $\Delta_{n,\lambda}$ has a filtration whose factors are exactly $\overline{\Delta}_{n,\lambda_{i}}$. Since $\lambda$ is singular, we can find a unique singular $\mu\in\operatorname{Irr}(G_{n+1})$ together with a nonzero homomorphism $\phi:\Delta_{n+1,\mu}\to\Delta_{n,\lambda}$ by Theorem 5.1. Therefore, there exists a nonzero homomorphism from $\Delta_{n+1,\mu}$ to $\overline{\Delta}_{n,\lambda_{i}}$ for a certain $i\in[r]$. Note that $\lambda_{i}$ must be singular. Otherwise, $\overline{\Delta}_{n,\lambda_{i}}$ is irreducible by Proposition 5.6. But this is impossible since the top of $\Delta_{n+1,\mu}$ is not isomorphic to $\overline{\Delta}_{n,\lambda_{i}}$.
By Lemma 5.7, the length of $\overline{\Delta}_{n,\lambda_{i}}$ is at most 2. Since the above map from $\Delta_{n+1,\mu}$ to $\overline{\Delta}_{n,\lambda_{i}}$ is clearly not surjective, its image is contained in the socle of $\overline{\Delta}_{n,\lambda_{i}}$. Consequently, one gets a nonzero homomorphism
$\overline{\phi}:\overline{\Delta}_{n+1,\mu}\longrightarrow\overline{\Delta}_{n,\lambda_{i}},$
which is the desired homomorphism for ${\mathscr{C}}={\mathrm{OA}},{\mathrm{CA}},{\mathrm{FA}}$ since $\lambda_{i}\cong\lambda$ by Proposition 5.4: there is a unique singular irreducible representation in $\operatorname{Irr}(G_{n})$.
For ${\mathscr{C}}={\mathrm{BA}}$ or $\mathrm{SA}$}, $\lambda_{i}$ may not be isomorphic to $\lambda$, but the other singular irreducible $\lambda^{\prime}$ twisted by a sign. In this case, we repeat the above argument for $\lambda^{\prime}$ to get a nonzero homomorphism
$\overline{\phi}^{\prime}:\overline{\Delta}_{n+1,\mu^{\prime}}\longrightarrow\overline{\Delta}_{n,\lambda^{\prime}_{i}},$
where $\lambda_{i}^{\prime}\in\operatorname{Irr}(G_{n})$ is singular and $\mu^{\prime}$ is not isomorphic to $\mu$. Then $\lambda_{i}^{\prime}$ is isomorphic to $\lambda$ and we get the desired map. Indeed, if this is not the case, then $\lambda_{i}^{\prime}$ is isomorphic to $\lambda^{\prime}$ as well. But then we get two nonzero homomorphisms $\overline{\phi}$ and $\overline{\phi}^{\prime}$ to $\overline{\Delta}_{n,\lambda^{\prime}}$, whose images land in distinct simple submodules of the socle. Consequently, the length of $\overline{\Delta}_{n,\lambda^{\prime}}$ is at least 3, contradicting Lemma 5.7. ∎
Combining these two lemmas, we get:
Proposition 5.9.
Let $k$ be a field. If $\lambda\in\operatorname{Irr}(G_{n})$ is singular, then $\overline{\Delta}_{n,\lambda}$ has length 2.
Proof.
Lemma 5.7 tells us that $\overline{\Delta}_{n,\lambda}$ has length at most 2, and Lemma 5.8 asserts that its length is at least 2. The conclusion follows. ∎
@@
Let $k$ be a field. Then the isomorphism classes of irreducible $\mathscr{C}$}-modules are parameterized by elements in the set
@@
    if $\lambda$ is regular, then the corresponding irreducible $\mathscr{C}$}-module is isomorphic to $\overline{\Delta}_{n,\lambda}$;
@@
    if $\lambda$ is singular, then the corresponding irreducible $\mathscr{C}$}-module is the top of $\overline{\Delta}_{n,\lambda}$.
@@
Let $k$ be a field. Then the Grothendieck group $K_{0}({\mathscr{C}})$ of the category of finitely generated $\mathscr{C}$}-modules has the following canonical decomposition:
@@
Next we show that the set $\{[\Delta_{n,\lambda}]\mid n\in\mathbb{Z}_{+},\,\lambda\in\operatorname{Irr}(G_{n})\}$ is a basis as well. By Proposition 5.6 and 5.9, if $\lambda\in\operatorname{Irr}(G_{n})$ is regular, then $\Delta_{n,\lambda}]=[L_{n,\lambda}]$; otherwise,
@@
by the proof of Proposition 5.9. Therefore, the transition matrix between $\{[L_{n,\lambda}]\}$ and $\{[\Delta_{n,\lambda}]\}$ is upper unitriangular. Since an upper unitriangular matrix (even of infinite size, as long as the degrees are bounded below) is invertible over $\mathbb{Z}$, the classes $\{[\Delta_{n,\lambda}]\}$ form an alternative basis for $K_{0}({\mathscr{C}})$. ∎
@@ 6. Further results for $\mathrm{CA}$}, $\mathrm{SA}$} and $\mathrm{FA}$}
In the case that $k$ is a field, the representation theory of $\mathrm{OA}$} and $\mathrm{BA}$} is very transparent. Thus in this section we only consider representations of the other three categories.
@@ 6.1. $\mathrm{CA}$} and $\mathrm{SA}$}
Since $d$ is either 1 or $2$, this computation tells us that most indecomposable standard modules are actually projective. More precisely, fix $n\geqslant 1$, and let
@@
$k^{d}\cong\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n})=\bigoplus_{i=1}^{r}\operatorname{Ext}_{k{\mathscr{C}}}^{1}((\Delta_{n+1,\mu_{i}})^{a_{i}},\,\Delta_{n})$
@@
$\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n})=\bigoplus_{i=1}^{r}\bigoplus_{j=1}^{s}\operatorname{Ext}_{k{\mathscr{C}}}^{1}((\Delta_{n+1,\mu_{j}})^{a_{j}},\,(\Delta_{n,\lambda_{i}})^{b_{i}})$
@@
$\Delta_{n,\lambda}(n+1)\cong k{\mathscr{C}}^{+}(n,n+1)\otimes_{kG_{n}}kG_{n}e_{\lambda},$
which has dimension $(n+1)\dim_{k}(kG_{n}e_{\lambda})$ since $k{\mathscr{C}}^{+}(n,n+1)$ is a right free $kG_{n}$-module of rank $n+1$. For the third term, we note that $\mathfrak{I}_{n+1}$ has a short exact sequence
@@
has dimension $(n+1)\dim_{k}(kG_{n}e_{\lambda})$. Consequently, $\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1},\,\Delta_{n,\lambda})$ has dimension 1, so it is nonzero as claimed.
Let $e_{\mu}\in kG_{n+1}$ be a primitive idempotent such that $kG_{n+1}e_{\mu}$ is the projective cover of $\mu$. Applying $\operatorname{Hom}_{k{\mathscr{C}}}(-,\,\Delta_{n})$ to the short exact sequence
@@
which has dimension $n\dim_{k}(kG_{n+1}e_{\mu})$ since $\Delta_{n}(n+1)\cong k{\mathscr{C}}^{+}(n,n+1)$ is a free left $kG_{n+1}$-module of rank $n$. Finally, $\mathfrak{I}_{n+1}e_{\mu}$ admits a filtration whose only contribution to $\operatorname{Hom}(-,\,\Delta_{n})$ comes from
$k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}k{\mathscr{C}}^{-}(n+1,n)e_{\mu}\cong k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}(kG_{n}\otimes_{k}kG_{n+1}e_{\mu})\cong(k{\mathscr{C}}\otimes_{k{\mathscr{C}}^{-}}kG_{n})\otimes_{k}kG_{n+1}e_{\mu}\cong\Delta_{n}\otimes_{k}kG_{n+1}e_{\mu},$
@@
$\operatorname{Hom}_{k{\mathscr{C}}}(\mathfrak{I}_{n+1}e_{\mu},\,\Delta_{n})\cong\operatorname{Hom}_{k{\mathscr{C}}}(\Delta_{n},\,\Delta_{n}\otimes_{k}kG_{n+1}e_{\mu})$
has dimension $n\dim_{k}(kG_{n+1}e_{\mu})$, so $\operatorname{Ext}_{k{\mathscr{C}}}^{1}(\Delta_{n+1,\mu},\,\Delta_{n})$ has dimension 1, which is nonzero as claimed.
For the category $\mathrm{SA}$}, the argument proceeds similarly with a refined look at the bimodule structure of the morphisms. For $n\geqslant 3$, $G_{n+1}=D_{2(n+1)}$ and $G_{n}=D_{2n}$ are dihedral, so $k{\mathscr{C}}^{-}(n+1,n)$ is not a free bimodule. Instead, since the action of $G_{n}\times G_{n+1}$ on ${\mathscr{C}}^{-}(n+1,n)$ is transitive and the stabilizer is a subgroup of order 2, one has
@@
$k{\mathscr{C}}^{-}(n+1,n)e_{\mu}\cong kG_{n}\otimes_{kC_{2}}kG_{n+1}e_{\mu},$
which has dimension $n\dim_{k}(kG_{n+1}e_{\mu})$ since $kG_{n}$ is a free right $kC_{2}$-module of rank $n$. Consequently, the dimension of $\operatorname{Hom}_{k{\mathscr{C}}}(\mathfrak{I}_{n+1}e_{\mu},\,\Delta_{n})$ remains $n\dim_{k}(kG_{n+1}e_{\mu})$, preserving the exact dimension balance required to show that $\operatorname{Ext}^{1}_{k{\mathscr{C}}}(\Delta_{n+1,\mu},\,\Delta_{n,\lambda})$ is one-dimensional. ∎